# Domain-Specific Translation with Open-Source Large Language Models: Resource-Oriented Analysis

### v2

@@ Abstract
In this work, we compare the domain-specific translation performance of open-source autoregressive decoder-only large language models (LLMs) with task-oriented machine translation (MT) models. Our experiments focus on the medical domain and cover four language pairs with varied resource availability: English-to-French, English-to-Portuguese, English-to-Swahili, and Swahili-to-English. Despite recent advancements, LLMs exhibit a clear gap in specialized translation quality compared to multilingual encoder-decoder MT models such as NLLB-200. In three out of four language directions in our study, NLLB-200 3.3B outperforms all LLMs in the size range of 8B parameters in medical translation. While fine-tuning LLMs such as Mistral and Llama improves their performance at medical translation, these models still fall short compared to fine-tuned NLLB-200 3.3B models. Our findings highlight the ongoing need for specialized MT models to achieve higher-quality domain-specific translation, especially in medium-resource and low-resource settings. As larger LLMs outperform their 8B variants, this also encourages pre-training domain-specific medium-sized LMs to improve quality and efficiency in specialized translation tasks.
@@ 1 Introduction
In recent years, there have been several advancements related to employing large language models (LLMs) in diverse natural language processing (NLP) tasks, including machine translation (MT). Autoregressive decoder-only models such as BLOOM (Le Scao et al.,, 2022), ChatGPT (Brown et al.,, 2020), Mistral (Jiang et al.,, 2023), and Llama (Touvron et al.,, 2023) –to mention but a few– have been welcomed by NLP researchers and practitioners as a new paradigm, due to their outstanding in-context learning feature. As the name suggests, in-context learning enables an LLM to leverage additional context during inference to enhance its output (Brown et al.,, 2020). When it comes to MT, researchers investigated incorporating all sorts of context, such as translation pairs similar to the new source text (Agrawal et al.,, 2023; Moslem et al., 2023a, ; Vilar et al.,, 2023), terminology (Moslem et al., 2023c, ), and dictionary words (Ghazvininejad et al.,, 2023). Usually prompting the model, i.e. providing instructions in a natural language, is sufficient without additional model fine-tuning. Nevertheless, several researchers investigated fine-tuning LLMs for MT tasks, for both sentence-level translation (Alves et al.,, 2023; Xu et al.,, 2023), and adaptive translation (Moslem et al., 2023b, ; Zhang et al.,, 2023; Stap et al.,, 2024).
@@
  • clear gaps when it comes to language support, whether for the generic domain (Robinson et al.,, 2023), or for specialized domains (Moslem et al., 2023a, ).
@@
This reopens the door to the question whether we still need conventional encoder-decoder MT models. In this work, we focus on domain-specific translation, and use NLLB-200 3.3B, both the baseline and fine-tuned models, as our main task-oriented benchmarks. NLLB-200 3.3B has demonstrated high quality comparable with strong commercial systems in diverse languages (Moslem et al., 2023a, ). The majority of decoder-only LLMs we explore are in the size range of 8B parameters, namely Mistral 7B (Jiang et al.,, 2023), Llama-3 8B, Llama-3.1 8B (Dubey et al.,, 2024), Gemma 7B (Mesnard et al.,, 2024), Qwen’s QwQ-32B-Preview, and Qwen-2.5 72B (Qwen et al.,, 2024). Moreover, we evaluate Llama-3 70B, and Mixtral 8x7B, which is a mixture of 7B experts model, as well as the massive Llama-3.1 405B and DeepSeek (DeepSeek-AI et al.,, 2024).
@@ 2.1 French and Portuguese
We gathered medical data for English-to-Portuguese (EN-PT) and English-to-French (EN-FR) translations from the OPUS repository, specifically the ELRC, EMEA (Tiedemann,, 2012), SciELO (Soares et al.,, 2018), and TICO-19 (Anastasopoulos et al.,, 2020) datasets. Following a combination of rule-based and semantic filtering techniques, we curated 2,905,418 segments for EN-PT and 303,948 segments for EN-FR. [^2] Data preparation scripts are available at: https://github.com/ymoslem/MT-Preparation/
@@ 2.2 Swahili
We collected medical and generic datasets for Swahili-to-English (SW-EN) and English-to-Swahili (EN-SW) models from OPUS (Tiedemann,, 2012). We used Global Voices and Wikimedia for the generic domain, and TICO-19 (Anastasopoulos et al.,, 2020) and Wikipedia articles for the medical domain. After applying rule-based and semantic filtering, we obtained 39,415 segments for the generic domain and 3,514 segments for the medical domain.
@@ 3 Experiments
In our experiments, we start by evaluating the baselines of various open-source models, including both autoregressive decoder-only LLMs, and the encoder-decoder NLLB-200 3.3B model, for both zero-shot and one-shot translation. Then, we fine-tune the models that show the best performance. In most cases, the LLMs chosen for fine-tuning are Mistral 7B and Llama-3 8B for their translation quality and efficiency. Following Moslem et al., 2023a ,  [^3] Code for the experiments: https://github.com/ymoslem/Adaptive-MT-LLM-Fine-tuning we fine-tune these LLMs using a combination of zero-shot and one-shot prompts in the medical domain. For the one-shot prompt experiments, we augment each translation pair with an additional translation example that is a fuzzy match to the source sentence. We extract fuzzy matches from a unique dataset which we refer to as the “context dataset” (cf. Section 2). Fuzzy matches are identified through calculating the semantic similarity between the embeddings of the source segments in the training datasets and those in the “context dataset”. The top match is selected for a one-shot prompt. Both the baseline and the fine-tuned models are then evaluated separately on zero-shot and one-shot prompts.
@@
Fuzzy match retrieval involves three steps: (i) creating the embeddings, (ii) indexing, and (iii) retrieval. For creating the embeddings, we use Sentence-Transformers (Reimers and Gurevych,, 2019) with Microsoft Multilingual-MiniLM-L12-H384. Additionally, we tested other embedding models, and reported the results in Appendix B. For indexing, we use Faiss (Johnson et al.,, 2019), a library for efficient similarity search and clustering of dense vectors, training an IndexIVFFlat index with IndexFlatL2 as a quantizer. Retrieval was done by similarity search with cosine similarity.
@@
In our experiments, we fine-tune a range of autoregressive decoder-only models, namely: Mistral 7B, Mixtral 7x8B, Llama-3 8B, and Gemma 7B. For efficiency, we employ QLoRA, a parameter-efficient fine-tuning technique. We fine-tune the models for one epoch, as in Moslem et al., 2023a , following the same configuration settings. As Llama-3 8B revealed the best results (cf. Section 4), we experimented with fine-tuning it with two sizes of data sets, (i) small [s+1], which is used for fine-tuning all the other models, and (ii) and medium [m+1]. The details of these datasets can be referred to in Section 2.
@@
Moreover, we evaluated the NLLB-200 3.3B model, which is a state-of-the-art sequence-to-sequence machine translation model, based on the encoder-decoder Transformer architecture. For NLLB-200 3.3B fine-tuning, we conducted 4 experiments: (i) sentence-level fine-tuning with the small dataset [s]; (ii) sentence-level fine-tuning with the medium dataset [m]; (iii) concatenation of a fuzzy match with the new sentence pair in the small dataset [s+1]; and (iv) concatenation of a fuzzy match with the new sentence pair in the medium dataset [m+1]. The concatenated fuzzy segment and new segment are separated by both the language code and a special separator (Bulte and Tezcan,, 2019; Moslem et al., 2023b, ). At the decoding time, for the two models that use fuzzy-match concatenation, we apply teacher forcing of the fuzzy target as a translation prefix so that the model completes the translation of the new source accordingly. [^4] Refer to the CTranslate2’s “target_prefix” option.
@@
For efficient inference, we use vLLM (Kwon et al.,, 2023) for translation with all autoregressive models, while we use CTranslate2 for NLLB-200 3.3B, which is not supported by vLLM. [^5] CTranslate2 supports some, but not all the autoregressive models in our experiments; hence, we decided to use vLLM for these models. Appendix C compares scores obtained from CTranslate2 and vLLM. Due to the massive size of Llama 3.1 405B and DeepSeek V3 685B, we had to use Activation-aware Weight Quantization (AWQ) (Lin et al.,, 2023), while other baseline models do not use quantization at inference time.
@@ 4 Evaluation and Discussion
In this section, we discuss the evaluation results of our experiments. We experimented with both pure prompting pretrained LLMs for the task of translation, and fine-tuning these models. In both set of experiments, we evaluate two scenarios: (i) sentence-level translation (zero-shot), and (ii) retrieval-augmented translation through incorporating one in-context example (fuzzy match) into translation prompts (one shot). To evaluate our systems, we calculated BLEU (Papineni et al.,, 2002), and chrF++ (Popović,, 2017), as implemented in the sacreBLEU library (Post,, 2018). [^6] https://github.com/mjpost/sacrebleu For semantic evaluation, we used COMET Rei et al., (2020). [^7] https://github.com/Unbabel/COMET, [^8] In particular, we used the “wmt20-comet-da” model. As COMET does not fully support Swahili, we used AfriCOMET-MTL (multitask learning) (Wang et al.,, 2024) evaluation model. [^9] https://hf.co/masakhane/africomet-mtl
@@ 4.1.1 Baseline Evaluation
For zero-shot translation, NLLB-200 3.3B, the encoder-decoder MT model, outperforms all the baseline models in the size range of 8B parameters, while Llama-3 8B and Llama-3.1 one-shot translation outperforms the models in the same size range. Although Llama-3.1 models are promoted as being “multilingual” (Dubey et al.,, 2024), Llama-3 8B and Llama 3.1 8B show comparable performance.
@@
BioMistral (Labrak et al.,, 2024) is an open-source LLM tailored for the biomedical domain, utilizing Mistral as its foundation model and further pre-trained on PubMed Central. However, in our experiments, Mistral outperforms BioMistral at both zero-shot and one-shot EN-FR medical translation. Finally, Gemma 7B has the worst performance among all the models we evaluated for zero-shot translation. However, for one-shot translation, Gemma slightly outperforms Mistral.
@@ 4.1.2 Fine-tuning Evaluation
After fine-tuning Mistral, we observe considerable performance gains in both zero-shot and one-shot translation for both French and Portuguese (cf. Table 1 and Table 2). Interestingly, while fine-tuning Mixtral 8x7B enhances its evaluation scores for zero-shot translation, it degrades its one-shot translation performance. Fine-tuning both Mistral and Mixtral 8x7B, zero-shot translation of Mixtral 8x7B still outperforms Mistral 7B, while we do not observe the same level of improvement for one-shot translation. Fine-tuning Gemma 7B considerably improves zero-shot translation, while degrades one-shot EN-FR translation. Nevertheless, except for BLEU, Mistral still outperforms Gemma after fine-tuning both models for both zero-shot and one-shot translation based on all other metrics. The fine-tuned Llama-3 8B for EN-FR medical translation outperforms Llama-3 70B baseline (i.e. without fine-tuning) for zero-shot translation, while the evaluation scores for one-shot translation are relatively comparable. For EN-PT medical translation, the larger model Llama-3 70B still outperforms the fine-tuned Llama-3 8B model.
@@
It is worth mentioning that NLLB-200 3.3B was not originally pre-trained with fuzzy-match concatenation. Hence, one-shot translation might have quality issues in some translations, alleviated by fine-tuning. In other words, pre-training an MT model with domain-specific fuzzy-match concatenation can be considered (Bulte and Tezcan,, 2019) if this feature is required later for fine-tuning and inference. Moreover, using quality estimation tools such as CometKiwi (Rei et al.,, 2022) or MetricX (Juraska et al.,, 2024) is recommended to select the best translation from zero-shot (sentence-level) and one-shot (fuzzy-match concatenated) translations. Nevertheless, it is important to evaluate the quality gains of fuzzy-match concatenation per se. Fuzzy-match concatenation diversifies and increases the data, which might explain improvements in zero-shot translation, but not one-shot translation. As in the case of Portuguese and Swahili (cf. Section 4.2), conventional sentence-level fine-tuning can be sufficient. Other data augmentation techniques can be considered, including monolingual data back-translation (Sennrich et al.,, 2016; Edunov et al.,, 2018) and domain-specific text generation (Moslem et al.,, 2022).
@@ 4.2 Swahili
For both the medical domain and the generic domain (cf. Table 3) of Swahili translation, Gemma and Llama-3 8B (especially after fine-tuning) are the top performers among autoregressive decoder-only models in the size range of 8B parameters. However, their performance still lags behind that of NLLB-200 3.3B. Nevertheless, Llama-3 70B evaluation scores are relatively comparable to those of NLLB-200 3.3B.
@@
Although Llama 3.1 8B is promoted as a multilingual model Dubey et al., (2024), Llama 3 8B outperforms it for EN-SW translation. Comparing model sizes, Llama 3 70B outperforms the smaller version of the model Llama 3 8B for both zero-shot and one-shot EN-SW translation.

### v4

@@ Abstract
In this work, we compare the domain-specific translation performance of open-source autoregressive decoder-only large language models (LLMs) with task-oriented machine translation (MT) models. Our experiments focus on the medical domain and cover four language directions with varied resource availability: English-to-French, English-to-Portuguese, English-to-Swahili, and Swahili-to-English. Despite recent advancements, LLMs demonstrate a significant quality gap in specialized translation compared to multilingual encoder-decoder MT models such as NLLB-200. Our results indicate that NLLB-200 3.3B outperforms all evaluated LLMs in the 7-8B parameter range across three out of the four language directions. While fine-tuning improves the performance of LLMs such as Mistral and Llama, these models still underperform compared to fine-tuned NLLB-200 3.3B models. Our findings highlight the ongoing need for specialized MT models to achieve high-quality domain-specific translation, especially in medium-resource and low-resource settings. Moreover, the superior performance of larger LLMs over their 8B variants suggests potential value in pre-training domain-specific medium-sized language models, employing targeted data selection and knowledge distillation approaches to enhance both quality and efficiency in specialized translation tasks.
@@ 1 Introduction
In recent years, there have been several advancements related to employing large language models (LLMs) in diverse natural language processing (NLP) tasks, including machine translation (MT). Autoregressive decoder-only models such as BLOOM (Le Scao et al., 2022), ChatGPT (Brown et al., 2020), Mistral (Jiang et al., 2023), and Llama (Touvron et al., 2023) –to mention but a few– have been welcomed by NLP researchers and practitioners as a new paradigm, due to their outstanding in-context learning feature. As the name suggests, in-context learning enables an LLM to leverage additional context during inference to enhance its output (Brown et al., 2020). When it comes to MT, researchers investigated incorporating all sorts of context, such as translation pairs similar to the new source text (Agrawal et al., 2023; Moslem et al., 2023a; Vilar et al., 2023), terminology (Moslem et al., 2023c), and dictionary words (Ghazvininejad et al., 2023). Usually prompting the model, i.e. providing instructions in a natural language, is sufficient without additional model fine-tuning. Nevertheless, several researchers investigated fine-tuning LLMs for MT tasks, for both sentence-level translation (Alves et al., 2023; Xu et al., 2023), and adaptive translation (Moslem et al., 2023b; Zhang et al., 2023; Stap et al., 2024).
@@
  • clear gaps when it comes to language support, whether for the generic domain (Robinson et al., 2023), or for specialized domains (Moslem et al., 2023a).
@@
This reopens the door to the question whether we still need conventional encoder-decoder MT models. In this work, we focus on domain-specific translation, and use NLLB-200 3.3B, both the baseline and fine-tuned models, as our main task-oriented benchmarks. NLLB-200 3.3B has demonstrated high quality comparable with strong commercial systems in diverse languages (Moslem et al., 2023a). The majority of decoder-only LLMs we explore are in the size range of 8B parameters, namely Mistral 7B (Jiang et al., 2023), Llama-3 8B, Llama-3.1 8B (Dubey et al., 2024), Gemma 7B (Mesnard et al., 2024), Qwen’s QwQ-32B-Preview, and Qwen-2.5 72B (Qwen et al., 2024). Moreover, we evaluate Llama-3 70B, and Mixtral 8x7B, which is a mixture of 7B experts model, as well as the massive Llama-3.1 405B and DeepSeek (DeepSeek-AI et al., 2024).
@@ 2.1 French and Portuguese
We gathered medical data for English-to-Portuguese (EN-PT) and English-to-French (EN-FR) translations from the OPUS repository, specifically the ELRC, EMEA (Tiedemann, 2012), SciELO (Soares et al., 2018), and TICO-19 (Anastasopoulos et al., 2020) datasets. Following a combination of rule-based and semantic filtering techniques, we curated 2,905,418 segments for EN-PT and 303,948 segments for EN-FR. [^2] Data preparation scripts are available at: https://github.com/ymoslem/MT-Preparation/
@@ 2.2 Swahili
We collected medical and generic datasets for Swahili-to-English (SW-EN) and English-to-Swahili (EN-SW) models from OPUS (Tiedemann, 2012). We used Global Voices and Wikimedia for the generic domain, and TICO-19 (Anastasopoulos et al., 2020) and Wikipedia articles for the medical domain. After applying rule-based and semantic filtering, we obtained 39,415 segments for the generic domain and 3,514 segments for the medical domain.
@@ 3 Experiments
In our experiments, we start by evaluating the baselines of various open-source models, including both autoregressive decoder-only LLMs, and the encoder-decoder NLLB-200 3.3B model, for both zero-shot and one-shot translation. Then, we fine-tune the models that show the best performance. In most cases, the LLMs chosen for fine-tuning are Mistral 7B and Llama-3 8B for their translation quality and efficiency. Following Moslem et al. (2023a),  [^3] Code for the experiments: https://github.com/ymoslem/Adaptive-MT-LLM-Fine-tuning we fine-tune these LLMs using a combination of zero-shot and one-shot prompts in the medical domain. For the one-shot prompt experiments, we augment each translation pair with an additional translation example that is a fuzzy match to the source sentence. We extract fuzzy matches from a unique dataset which we refer to as the “context dataset” (cf. Section 2). Fuzzy matches are identified through calculating the semantic similarity between the embeddings of the source segments in the training datasets and those in the “context dataset”. The top match is selected for a one-shot prompt. Both the baseline and the fine-tuned models are then evaluated separately on zero-shot and one-shot prompts.
@@
Fuzzy match retrieval involves three steps: (i) creating the embeddings, (ii) indexing, and (iii) retrieval. For creating the embeddings, we use Sentence-Transformers (Reimers and Gurevych, 2019) with Microsoft Multilingual-MiniLM-L12-H384. Additionally, we tested other embedding models, and reported the results in Appendix B. For indexing, we use Faiss (Johnson et al., 2019), a library for efficient similarity search and clustering of dense vectors, training an IndexIVFFlat index with IndexFlatL2 as a quantizer. Retrieval was done by similarity search with cosine similarity.
@@
In our experiments, we fine-tune a range of autoregressive decoder-only models, namely: Mistral 7B, Mixtral 7x8B, Llama-3 8B, and Gemma 7B. For efficiency, we employ QLoRA, a parameter-efficient fine-tuning technique. We fine-tune the models for one epoch, as in Moslem et al. (2023a), following the same configuration settings. As Llama-3 8B revealed the best results (cf. Section 4), we experimented with fine-tuning it with two sizes of data sets, (i) small [s+1], which is used for fine-tuning all the other models, and (ii) and medium [m+1]. The details of these datasets can be referred to in Section 2.
@@
Moreover, we evaluated the NLLB-200 3.3B model, which is a state-of-the-art sequence-to-sequence machine translation model, based on the encoder-decoder Transformer architecture. For NLLB-200 3.3B fine-tuning, we conducted 4 experiments: (i) sentence-level fine-tuning with the small dataset [s]; (ii) sentence-level fine-tuning with the medium dataset [m]; (iii) concatenation of a fuzzy match with the new sentence pair in the small dataset [s+1]; and (iv) concatenation of a fuzzy match with the new sentence pair in the medium dataset [m+1]. The concatenated fuzzy segment and new segment are separated by both the language code and a special separator (Bulte and Tezcan, 2019; Moslem et al., 2023b). At the decoding time, for the two models that use fuzzy-match concatenation, we apply teacher forcing of the fuzzy target as a translation prefix so that the model completes the translation of the new source accordingly. [^4] Refer to the CTranslate2’s “target_prefix” option.
@@
For efficient inference, we use vLLM (Kwon et al., 2023) for translation with all autoregressive models, while we use CTranslate2 for NLLB-200 3.3B, which is not supported by vLLM. [^5] CTranslate2 supports some, but not all the autoregressive models in our experiments; hence, we decided to use vLLM for these models. Appendix C compares scores obtained from CTranslate2 and vLLM. Due to the massive size of Llama 3.1 405B and DeepSeek V3 685B, we had to use Activation-aware Weight Quantization (AWQ) (Lin et al., 2023), while other baseline models do not use quantization at inference time.
@@ 4 Evaluation and Discussion
In this section, we discuss the evaluation results of our experiments. We experimented with both pure prompting pretrained LLMs for the task of translation, and fine-tuning these models. In both set of experiments, we evaluate two scenarios: (i) sentence-level translation (zero-shot), and (ii) retrieval-augmented translation through incorporating one in-context example (fuzzy match) into translation prompts (one shot). To evaluate our systems, we calculated BLEU (Papineni et al., 2002), and chrF++ (Popović, 2017), as implemented in the sacreBLEU library (Post, 2018). [^6] https://github.com/mjpost/sacrebleu For semantic evaluation, we used COMET Rei et al. (2020). [^7] https://github.com/Unbabel/COMET, [^8] In particular, we used the “wmt20-comet-da” model. As COMET does not fully support Swahili, we used AfriCOMET-MTL (multitask learning) (Wang et al., 2024) evaluation model. [^9] https://hf.co/masakhane/africomet-mtl
@@ 4.1.1 Baseline Evaluation
For zero-shot translation, NLLB-200 3.3B, the encoder-decoder MT model, outperforms all the baseline models in the size range of 8B parameters, while Llama-3 8B and Llama-3.1 one-shot translation outperforms the models in the same size range. Although Llama-3.1 models are promoted as being “multilingual” (Dubey et al., 2024), Llama-3 8B and Llama 3.1 8B show comparable performance.
@@
BioMistral (Labrak et al., 2024) is an open-source LLM tailored for the biomedical domain, utilizing Mistral as its foundation model and further pre-trained on PubMed Central. However, in our experiments, Mistral outperforms BioMistral at both zero-shot and one-shot EN-FR medical translation. Finally, Gemma 7B has the worst performance among all the models we evaluated for zero-shot translation. However, for one-shot translation, Gemma slightly outperforms Mistral.
@@ 4.1.2 Fine-tuning Evaluation
After fine-tuning Mistral, we observe considerable performance gains in both zero-shot and one-shot translation for both French and Portuguese (cf. Table 1 and Table 2). Interestingly, while fine-tuning Mixtral 8x7B enhances its performance for zero-shot translation, it degrades its one-shot translation performance. Fine-tuning both Mistral and Mixtral 8x7B, zero-shot translation of Mixtral 8x7B still outperforms Mistral 7B, while we do not observe the same level of improvement for one-shot translation. Fine-tuning Gemma 7B considerably improves zero-shot translation, while degrades one-shot EN-FR translation. Nevertheless, except for BLEU, Mistral still outperforms Gemma after fine-tuning both models for both zero-shot and one-shot translation based on all other metrics. The fine-tuned Llama-3 8B for EN-FR translation outperforms Llama-3 70B baseline (i.e. without fine-tuning) for zero-shot translation, while the evaluation scores for one-shot translation are relatively comparable. For EN-PT medical translation, the larger model Llama-3 70B still outperforms the fine-tuned Llama-3 8B model.
@@
It is worth mentioning that NLLB-200 3.3B was not originally pre-trained with fuzzy-match concatenation. Hence, one-shot translation might have quality issues in some translations, alleviated by fine-tuning. In other words, pre-training an encoder-decoder MT model with domain-specific fuzzy-match concatenation can be considered (Bulte and Tezcan, 2019) if this feature is required later for fine-tuning and inference. Moreover, using quality estimation tools such as CometKiwi (Rei et al., 2022) or MetricX (Juraska et al., 2024) is recommended to select the best translation from zero-shot (sentence-level) and one-shot (fuzzy-match concatenated) translations. Nevertheless, it is important to evaluate the quality gains of fuzzy-match concatenation per se. Fuzzy-match concatenation diversifies and increases the data, which might explain improvements in zero-shot translation, but not one-shot translation. As in the case of Portuguese and Swahili (cf. Section 4.2), conventional sentence-level fine-tuning can be sufficient. Other data augmentation techniques can be considered, including monolingual data back-translation (Sennrich et al., 2016; Edunov et al., 2018) and domain-specific text generation (Moslem et al., 2022).
@@ 4.2 Swahili
As illustrated by Table 3, in both the medical and generic domains of Swahili translation, Gemma and Llama-3 8B (especially after fine-tuning) are the top performers among autoregressive decoder-only LLMs in the size range of 8B parameters. However, NLLB-200 3.3B outperforms these models. Nevertheless, Llama-3 70B evaluation scores are relatively comparable to those of NLLB-200.
@@
Although Llama 3.1 8B is promoted as a multilingual model Dubey et al. (2024), Llama 3 8B outperforms it for EN-SW translation. Comparing model sizes, Llama 3 70B outperforms the smaller version of the model Llama 3 8B for both zero-shot and one-shot EN-SW translation.